OpenStudy (anonymous):

The sum of two consecutive even integers is 226. What is the larger integer?

OpenStudy (anonymous):

114

hero (hero):

When summing two even consecutive integers, you will have the general form x + (x+2) = n where n = the given integer.

hero (hero):

x + 2 will be the larger integer

hero (hero):

In this case, we have: x + (x+ 2) = 226 2x + 2 = 226 2x = 224 x = 224/2 x = 112 (the value of the smaller integer) x + 2 = 114 (the value of the larger integer)

OpenStudy (anonymous):

Thanks BUNCHES i never knew how to find that.

hero (hero):

yw